OpenStudy (anonymous):

To solve by completing the square, what value should you add to each side of the equation? x^2+10x=-3

jimthompson5910 (jim_thompson5910):

Edutopia you add 25 to both sides

jimthompson5910 (jim_thompson5910):

no it's simply asking what you add to both sides, which is 25

OpenStudy (anonymous):

ok ill delete my answer

OpenStudy (anonymous):

You have to make the equation equal to 0 before you can complete the square. x^2+10x=-3 x^2 + 10x + 3 = -3 +3 x^2 + 10x + 3 = 0 Then you take the second value (b), divide it by 2, and square it. c = (b/2)^2 c = (10/2)^2 c = (5)^2 c = 25 So then you add to the equation 25: (x^2 + 10x +25) - 25 + 3 = 0 <--- Don't forget you're both adding and subtracting the 25, to make them cancel each other out. Now take the terms in the brackets, and form a perfect square equation with them, and clean up the outside values (-25 and 3) (x+5)^2 + 22 = 0 And voila. There's your final equation.

OpenStudy (anonymous):

how would you get an integer or a simplified fraction out of this for an answer or is it not possible?

jimthompson5910 (jim_thompson5910):

dubbykinz you definitely went above and beyond of what it's asking (which is not a bad thing since it's helpful to know this stuff) but all they wanted to know is what to add to both sides, which in this case is 25

OpenStudy (anonymous):

I know, but a full explanation helps in understanding. :p no point in just saying "add 25" when you don't know what it's even for. Anyway I found the question weird because you're not really "adding" 25 to both sides of the equation. You're adding and subtracting 25 from the SAME side.

jimthompson5910 (jim_thompson5910):

well I mean you of course show how to get 25 by halving 10 to get 5, then squaring it to get 25 but the other stuff is extra it's definitely helpful though